### v1

@@ 4 Hydra: a novel method to mitigate CF
We observed that intraCF contributes more to the phenomenon of forgetting in SFL than the model drifting of part-1, with the last layers playing an important role. Methods in related work concerning CL or FL often utilize multiple blocks of higher layers, where each one of them is allocated to a different task, e.g., gurbuz2022nispa; kim2022multi; chen2023flexibility; hemati2025continual. In such methods, often called multi-head, the lower layers are shared among all tasks, while there exist multiple higher layers. We notice that the design of SFL naturally aligns with the multi-head architecture, as the higher layers are handled only by the server. Leveraging this together with the benefit of averaging under heterogeneous data (as shown by the accuracy and PG of FL in Fig. 3), we propose the first multi-head solution for SFL, named Hydra, to tackle CF in SFL. Hydra derives from our insights in Sec. 3.
Design Details. Fig. 4 presents the main elements of Hydra’s design and workflow. In detail, part-1 is trained in the same way as before (solely handled by the clients), while part-2 (managed by the server) is split and composed of: (i) one part-2a, shared among all clients and updated sequentially, and (ii) $G$ different versions of part-2b (heads), which are updated in parallel and aggregated (through FedAvg) at the end of each round. Further, for every version of part-2b, the data corresponding to a group of clients is assigned based on the highly represented labels in their data [^3] For this reason, we use the terms of “versions of part-2b”, “heads”, and “groups” interchangeably.. The server performs a mapping, i.e., assigns the different outputs (activations) of part-2a to different versions of part-2b during forward propagation. A natural choice for the number of part-2b’s versions is the number of labels, i.e., $G=L$, following the multi-head approaches in CL. Yet, $G$ can be tuned according to the server’s memory capacity to reduce overhead, as we discuss below.
@@
The groups are defined at the beginning of the training to cluster clients with similar data distributions, reducing large gradients that could overwrite previous updates. Homogeneity within a group ensures minimal weight changes, effectively addressing CF. This grouping requires each client to send information on the distribution of labels in its local data (in a form of a vector of length $L$) before the training starts. The grouping algorithm is based on resource allocation problems and aims at grouping clients with similar label distributions while ensuring that repartition of clients among the groups is balanced across all groups. In detail, it decides the assignments of clients to groups in a greedy way: for each group, the algorithm finds the client with the largest number of samples from the corresponding label and repeats this process until all clients have been assigned to exactly one group. Finally, we note that it iterates through the groups and assigns to them clients that are not assigned yet to any group. By doing so, and not iterating through the clients instead, the repartition of clients into the groups is balanced. We note that the same mapping is then used inversely by the server during backward propagation.
Unlike the existing multi-head approaches in CL, whose final trained model contains multiple heads, e.g., kim2022multi, Hydra’s final model contains a single head derived from aggregating all the part-2b’s versions. This is a key design feature that renders the multi-head approach more suitable for SFL. In fact, our evaluation (Sec. 5), showcases that aggregated heads excel in performance when compared to typical multi-head architectures. To the best of our knowledge, Hydra is the first attempt to fine-tune the multi-head technique for SFL. Further, another important design decision is the length of part-2b. To this end, we thoroughly study the choice of part-2b’s length in Sec. 5. Finally, Hydra differentiates from the USplit definition vepakomma2018split, in which the last (3rd) model part is processed at the clients and aggregated, since Hydra groups according to labels, which makes the last part less prone to model drifting.
@@
Table 2: Performance gap and global accuracy achieved by MobileNet with $\phi=10$.
Performance Gap $\downarrow$ Global Accuracy $\uparrow$
Dataset Policy Sharding Dir.α=0.3 Dir.α=0.1 Sharding Dir.α=0.3 Dir.α=0.1
CIFAR10 SFL+Hydra $\mathbf{10}$ $\mathbf{\pm 0.4}$ $\mathbf{21}$ $\mathbf{\pm 1}$ $\mathbf{32}$ $\mathbf{\pm 0.8}$ $\mathbf{72}$ $\mathbf{\pm 0.4}$ $\mathbf{59}$ $\mathbf{\pm 0.4}$ $\mathbf{49}$ $\mathbf{\pm 0.2}$
IID:(10, 80) SFL $36$ $\pm 1$ $43.7$ $\pm 0.5$ $45.5$ $\pm 1$ $50$ $\pm 0.4$ $50$ $\pm 0.5$ $40$ $\pm 0.6$
TinyImageNet SFL+Hydra $\mathbf{49}$ $\mathbf{\pm 0.6}$ $\mathbf{44}$ $\mathbf{\pm 0.6}$ $\mathbf{44.5}$ $\mathbf{\pm 1}$ $\mathbf{34}$ $\mathbf{\pm 0.1}$ $\mathbf{45}$ $\mathbf{\pm 0.1}$ $\mathbf{41}$ $\mathbf{\pm 0.3}$
IID:(40, 50) SFL $62$ $\pm 1.6$ $46$ $\pm 1$ $63$ $\pm 0.7$ $27$ $\pm 0.3$ $\mathbf{45}$ $\mathbf{\pm 0.3}$ $31$ $\pm 1$
Hydra’s Overhead. In terms of computing cost, there is a one-time $O(G\cdot C)$ cost for running the grouping algorithm at the beginning of training to assign clients into the different part-2b’s versions. During the training of part-2, there is no additional overhead since only one part-2b (head) is updated per batch update. Regarding the aggregation of parts-2b at the end of every round, this can be implemented locally at the server and run in parallel with the aggregation of part-1. In terms of memory overhead, this depends mainly on the number $G$ (heads) and the length of part-2b, but also on the implementation specifics. In practice, when part-2b consists of only the last 2 layers, the memory overhead can be as little as $G\cdot 6$MB, or $G\cdot 10$MB when part-2b consists of the last 6 layers. This estimation concerns Mobilenet and CIFAR-10 with a batch size of $64$ and is based on the memory needed for storing the model weights and the intermediate activations/gradients. In fact, our evaluation in Sec. 5 encourages smaller part-2b since Hydra performs better when compared to the case with a larger part-2b. Similarly, we show that Hydra leads to reduced CF even with a small $G$. Finally, Hydra induces a negligible communication overhead at the beginning of the training where each client sends a vector containing the information on its label distribution. However, we stress that Hydra induces no computational overhead on the clients’ side, since it only concerns part-2 of the model.
@@ 5 Experiments
In this section, we evaluate the performance of Hydra, while additional experiments can be found in the Appendix. Fig. 5 presents the results from SFL with and without Hydra (i.e., SFL+Hydra and SFL, respectively). It contains multiple aspects: processing order of part-2a at the server, datasets, ML models, and selection of part-2’s length.
Processing Order of Part-2a. We first focus on the case where the part-2b in Hydra is small and, in particular, its length is only 2 layers, i.e., the second cut is located at layer $24$ and $33$ for MobileNet and ResNet101, respectively. We notice that Hydra (SFL+Hydra) under cyclic order (with any $\phi$) significantly outperforms the baseline (SFL). In detail, by employing Hydra, the labels’ performance gap decreases by up to $64.5\%$ and $75.4\%$ for MobileNet and ResNet101 on CIFAR-10, respectively, while, at the same time, the global accuracy increases by $58.9\%$ and $112.2\%$, respectively, approaching the accuracy under IID data. We notice that Hydra has a robust performance regardless of the client scale ($\phi$), unlike the baseline SFL. Looking at the results for the random order, we see that Hydra reduces PG by up to $73\%$, thus providing a stable training even with random order.
Length of Part-2b in Hydra. We focus now on Hydra with small and big lengths of part-2b (e.g., for the big part-2b, the second cut layer is at layers $20$ and $30$ for MobileNet and ResNet101, respectively) in Fig. 5. In general, for MobileNet, the two cases perform similarly, i.e, Hydra with smaller part-2b achieves slightly better PG, while Hydra with bigger part-2b has slightly better accuracy, indicating that MobileNet is not affected by part-2b’s length. However, ResNet101, a more complex model, seems to be more sensitive to the size of part-2b since Hydra with smaller part-2b always outperforms the case of bigger part-2b, in terms of both accuracy and PG, a side effect of model drifting. Therefore, a smaller length for part-2b seems to be the best choice (and more lightweight).
Types of Data Partitioning. Table 2 presents the evaluation results of Hydra in three additional types of data partitioning, namely, Dirichlet, with $\alpha=0.1$ and $0.3$, and Sharding method with two dominant labels. The results demonstrate that Hydra improves over the baseline (SFL) in all the tested data partitioning methods (5 in total, considering also the $60\%$- and $80\%$-DL), even in Dir.-$\alpha_{0.1}$, which results in highly non-IID data.
Table 3: Comparison of Hydra with state-of-the-art methods, in MobileNet with CIFAR-10 and $80\%$-DL.
@@
Table 4: Ablation study on the number of Hydra’s heads ($G$), with MobileNet with CIFAR-10 and $80\%$-DL.
Policy N Perf. Gap $\downarrow$ Gl. Accuracy $\uparrow$
heads Cyclic $\phi_{=10}$ random Cyclic $\phi_{=10}$ random
@@
SFL+Hydra $10$ $\mathbf{15.4}$$\mathbf{\pm 0.6}$ $\mathbf{13.4}$$\mathbf{\pm 1}$ $\mathbf{65}$$\mathbf{\pm 0.1}$ $\mathbf{66.5}$$\mathbf{\pm 0.1}$
@@
Comparison with State-of-the-Art Methods. In Table 3, we evaluate Hydra (i.e., SFL+Hydra) using a random processing order (i.e., capturing a more generic scenario) and MobileNet versus other methods, i.e., SplitFedV1/FL thapa2022splitfed, SplitFedV3 madaan2022vulnerability, Multihead chen2023flexibility, MergeSFL liao2023mergesfl, and SplitNN vepakomma2018split. We stress that these methods do not follow a specific processing order, as all model parts are trained simultaneously. Table 3 shows that Hydra significantly outperforms the state-of-the-art. Specifically, unlike Hydra, which consistently excels in all metrics (accuracy, PG, and BW), methods in the literature fail to improve on all metrics (indicating an unstable training). More importantly, Hydra outperforms the other multi-head approaches (i.e., SplitFedV3, MultiHead). This highlights that Hydra’s aggregation of the heads (in contrast to the aforementioned methods) is a key feature over the classic multi-head approaches. Finally, MergeSFL achieves global accuracy comparable to the one of Hydra. However, a significant drawback of MergeSFL lies in its implementation, which requires synchronization among clients and more frequent updates (due to micro-batches).
Closing the Labels’ Performance Gap. In order to better understand Hydra’s performance, we dive into the per-position accuracy (i.e., the accuracy of the labels that are dominant in clients that are processed at a certain position, see Sec. 3.1). In Fig. 6, we see that Hydra effectively reduces the labels’ performance gap and stabilizes the training, even in a highly heterogeneous scenario, i.e., $80\%$-DL. In detail, unlike the baseline (SFL), there is no significant gap among the different positions. We note that the peaks and troughs in the subplot on the right are a result of the processing order under non-IID data. Finally, the global accuracy improves over SFL, i.e., $44\%$ instead of $28\%$, after 100 rounds.
Ablation Study on the Number of Heads $\mathbf{G}$. As discussed in Sec. 4, Hydra’s memory overhead can be adjusted by reducing the size of part-2b and/or $G$. For the former, we showed that a smaller part-2b performs better. Table 4 reports the performance of Hydra for different values of $G$ for CIFAR-10. For CIFAR-10, we define $G$ superclasses based on the semantic relationships among labels, similar to bai2021clustering. We see that Hydra performs best when $G=L$ but still outperforms the baseline, i.e., SFL, in all metrics, even for a smaller number of heads. These results highlight an interesting trade-off between memory overhead and performance.
@@ 6 Conclusion and Discussion
Limitations and Future Work. Despite the absence of theoretical claims in this work, our empirical analysis of CF in SFL paves the way towards a much-needed theoretical/convergence analysis of SFL with respect to key parameters such as the cut layer and the processing order at the server. Such an analysis could also be the stepping stone towards establishing theoretical guarantees for Hydra. Further, an interesting direction for future work is to incorporate client selection policies into the analysis of CF, but also in Hydra, similar to related work in FL, e.g., lee2022preservation. Finally, our ablation study on the number of Hydra’s heads revealed that semantic relationships among labels may play an important role in CF and in SFL’s training, in general. While this aspect was not the focus of this work, we believe that it should be studied in the future, as it has been done for continual learning ramasesh2020anatomy.
Acknowledgements
This work was funded by the Horizon MSCA Postdoctoral Fellowship OPALS (grant agreement 101210495), the Spanish Ministry of Economic Affairs and Digital Transformation and the European Union-NextGenerationEU through the project 6G-RIEMANN (TSI-063000-2021-147) and MAP-6G (TSI-063000-2021-63), and the Smart Networks and Services Joint Undertaking (SNS JU) under the European Union’s Horizon Europe research and innovation programme through CONFIDENTIAL6G EU project (Grant Agreement No 101096435). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.
Overview of Appendix
  • Appendix A: discusses design details of the proposed approach, Hydra, i.e., the proposed algorithm, and the grouping algorithm. Hydra is first presented in Section 4.
  • Appendix B: presents the overview of several variations of Split Federated Learning, and one Federated Learning approach that targets to tackle Catastrophic Forgetting. Note that these methods are used as baseline algorithms during the numerical evaluation.
  • Appendix C: extends Related work and contains background information regarding weight regularization methods used in Continual Learning.
  • Appendix D: presents the details of the experimental set-up used throughout the paper.
  • Appendix E: explains the types of data partitions used in this paper to simulate a setup of data heterogeneity.
  • Appendix F: contains details for the processing orders studied in the paper, e.g., cyclic order, as well as an additional one, the cyclic-and-reverse order.
  • Appendix G: presents additional experiments of the systematic analysis studied in Section 3.
  • Appendix H: presents additional results complementing the numerical evaluation of Section 5.

### v2

@@ 4 Hydra: A Novel Method to Mitigate CF
We observed that intraCF contributes more to the phenomenon of forgetting in SFL than the model drifting of part-1, with the last layers playing an important role. Methods in related work concerning CL or FL often utilize multiple blocks of higher layers, where each one of them is allocated to a different task, e.g., (Gurbuz and Dovrolis, 2022; Kim et al., 2022; Chen et al., 2023; Hemati et al., 2025). In such methods, often called multi-head, the lower layers are shared among all tasks, while there exist multiple higher layers. We notice that the design of SFL naturally aligns with the multi-head architecture, as the higher layers are handled only by the server. Leveraging this together with the benefit of averaging under heterogeneous data (as shown by the accuracy and PG of FL in Fig. 3), we propose the first multi-head solution for SFL, named Hydra, to tackle CF in SFL. Hydra derives from our insights in Sec. 3.
Design Details. Hydra extends SFL by modifying its workflow at the level of the server. Fig. 4 presents Hydra’s design and workflow. In detail, part-1 is trained in the same way as before (solely handled by the clients), while part-2 (managed by the server) is split and composed of: (i) a unique part-2a, shared among all clients and updated sequentially, and (ii) $G$ multiple versions of part-2b (heads), which are updated in parallel and aggregated (through FedAvg) at the end of each round. Each head corresponds to a label or a superclass (i.e., multiple labels). Further, the data corresponding to a group of clients is assigned to a specific head based on the highly represented labels in their data. This is done through a mapping that the server performs, where it assigns the different outputs (activations) of part-2a to the different heads during forward propagation. A natural choice for the number of heads is the number of labels, i.e., $G=L$, following the multi-head approaches in CL. However, $G$ can be tuned according to the server’s memory capacity to reduce overhead, as we discuss below.
@@
The group of clients corresponding to each head is defined at the beginning of the training to cluster clients with similar data distributions, reducing large gradients that could overwrite previous updates. Homogeneity within a group ensures minimal weight changes, effectively addressing CF. This grouping requires each client to send information on the distribution of labels in its local data (as a vector of length $L$) before the training starts. The grouping algorithm is based on resource allocation problems and aims at grouping clients with similar label distributions while ensuring that the repartition of clients among the groups is balanced across all groups. In detail, it assigns clients to groups in a greedy way: for each group, the algorithm finds the client with the largest number of samples from the corresponding label and repeats this process until all clients have been assigned to exactly one group. Finally, by iterating through the groups instead of the clients, the repartition of clients into the groups is balanced. Note that the same mapping is then used inversely by the server during backward propagation.
Unlike the existing multi-head approaches in CL, whose final trained model contains multiple heads, e.g., Kim et al. (2022), Hydra’s final model contains a single head derived from aggregating all the different heads, making multi-head more suitable for SFL. In fact, our evaluation (Sec. 5) shows that aggregated heads excel in performance when compared to typical multi-head architectures. To the best of our knowledge, Hydra is the first attempt to fine-tune the multi-head technique for SFL. Also, other important design decisions of Hydra include the processing order of clients’ activations through part-2a and the length of heads (part-2b). To this end, we thoroughly study the impact of these design choices in Sec. 5. Finally, we note that Hydra differentiates from the SL variant USplit Vepakomma et al. (2018), in which the model is split into 3 parts, and the last (3rd) model part is processed at the clients and aggregated (at the end of each round). Indeed, in Hydra, the number $G$ of heads is bounded by $L$, making it less prone to model drifting w.r.t. USplit, where $C$ model parts are aggregated.
@@
Global Accuracy $\uparrow$ Performance Gap $\downarrow$
Dataset Policy Sharding Dirichlet ($\alpha=0.3$) Dirichlet ($\alpha=0.1$) Sharding Dirichlet ($\alpha=0.3$) Dirichlet ($\alpha=0.1$)
CIFAR10 SFL+Hydra $\mathbf{72}$ $\mathbf{\pm 0.4}$ $\mathbf{59}$ $\mathbf{\pm 0.4}$ $\mathbf{49}$ $\mathbf{\pm 0.2}$ $\mathbf{10}$ $\mathbf{\pm 0.4}$ $\mathbf{21}$ $\mathbf{\pm 1}$ $\mathbf{32}$ $\mathbf{\pm 0.8}$
IID:(80, 10) SFL $50$ $\pm 0.4$ $50$ $\pm 0.5$ $40$ $\pm 0.6$ $36$ $\pm 1$ $43.7$ $\pm 0.5$ $45.5$ $\pm 1$
TinyImageNet SFL+Hydra $\mathbf{34}$ $\mathbf{\pm 0.1}$ $\mathbf{45}$ $\mathbf{\pm 0.1}$ $\mathbf{41}$ $\mathbf{\pm 0.3}$ $\mathbf{49}$ $\mathbf{\pm 0.6}$ $\mathbf{44}$ $\mathbf{\pm 0.6}$ $\mathbf{44.5}$ $\mathbf{\pm 1}$
IID:(50, 40) SFL $27$ $\pm 0.3$ $\mathbf{45}$ $\mathbf{\pm 0.3}$ $31$ $\pm 1$ $62$ $\pm 1.6$ $46$ $\pm 1$ $63$ $\pm 0.7$
Table 2: Global accuracy and performance gap achieved by MobileNet with $\phi=10$ under different data partitions.
Hydra’s Overhead. In terms of computing cost, there is a one-time $O(G\cdot C)$ cost for running the grouping algorithm at the beginning of training to assign clients into the different heads. During the training of part-2, there is no additional overhead since only one head is updated per batch update. Regarding the aggregation of heads at the end of every round, this can be implemented locally at the server and run in parallel with the aggregation of part-1. We stress that Hydra induces no computational overhead on the clients’ side, since it only concerns part-2 of the model (at the server).
In terms of memory overhead, this depends mainly on the number $G$ and the length of the heads, but also on the implementation specifics. In practice, when the heads consist of only the last 2 layers, the memory overhead can be as little as $G\cdot 6$MB, or $G\cdot 10$MB when the heads consist of the last 6 layers. This estimation concerns Mobilenet and CIFAR-10 with a batch size of $64$, and is based on the memory needed for storing the model weights and the intermediate activations/gradients. In fact, our evaluation in Sec. 5 encourages small heads since Hydra performs better when compared to the case with larger heads. Similarly, we show that Hydra leads to reduced CF when compared to vanilla SFL, even with a small $G$. Finally, Hydra induces a negligible communication overhead at the beginning of the training, where each client sends a vector containing the information on its label distribution (that is needed for the mapping algorithm).
@@ 5 Experiments
In this section, we thoroughly evaluate the performance of the proposed method Hydra. Fig. 5 presents the results from SFL with and without Hydra (denoted by SFL+Hydra and SFL, respectively). These results concern multiple aspects: the processing order of part-2a at the server, datasets, ML models, and the selection of the heads’ length.
Processing Order of Part-2a. We first focus on the case where Hydra’s heads are small, and, in particular, their length is only 2 layers. This means that the second cut layer (i.e., the last layer of part-2a) is located at layer $24$ and $33$ for MobileNet and ResNet101, respectively. We notice that Hydra (SFL+Hydra) under cyclic order (with any $\phi$) significantly outperforms the baseline (SFL). In detail, by employing Hydra, the labels’ performance gap decreases by up to $64.5\%$ and $75.4\%$ for MobileNet and ResNet101 on CIFAR-10, respectively, while, at the same time, the global accuracy increases by $58.9\%$ and $112.2\%$, respectively, approaching the accuracy under IID data. Furthermore, Hydra has a robust performance regardless of the client scale ($\phi$), unlike the baseline SFL. From the results for the random order, we observe that Hydra reduces PG by up to $73\%$, thus providing a stable training even under a random order.
Length of Heads (Part-2b) in Hydra. We focus now (in Fig. 5) on Hydra with small and large lengths of heads (e.g., for the large heads, the second cut layer is at layers $20$ and $30$ for MobileNet and ResNet101, respectively). In general, for MobileNet, the two cases perform similarly, i.e, Hydra with small heads achieves slightly better PG, while Hydra with larger heads has slightly better accuracy, indicating that MobileNet is not affected by the heads’ length. However, ResNet101, a more complex model, seems to be more sensitive to the length of heads since Hydra with small heads always outperforms the case of larger heads, in terms of both accuracy and PG, a side effect of model drifting. Therefore, a small length of heads seems to be the best design choice not only in terms of memory overhead (as discussed in Sec. 4), but also in terms of performance.
Types of Data Partitioning. In Sec. 4, Hydra was defined for any data partitioning. While Fig. 5 presented evaluation results of Hydra for the DL partition, Table 2 presents results in three additional types of data partitioning, namely Dirichlet with $\alpha=0.1$ and $0.3$ (where lower $\alpha$ indicates higher heterogeneity), and Sharding method with two dominant labels. The results demonstrate that Hydra improves over the baseline (SFL) in all the tested data partitioning methods.
@@
Table 3: Comparison of Hydra with state-of-the-art methods in MobileNet with CIFAR-10 and $80\%$-DL.
Policy N Gl. Accuracy $\uparrow$ Perf. Gap $\downarrow$
heads cyclic random cyclic random
@@
SFL+Hydra $10$ $\mathbf{65}$$\mathbf{\pm 0.1}$ $\mathbf{66.5}$$\mathbf{\pm 0.1}$ $\mathbf{15.4}$$\mathbf{\pm 0.6}$ $\mathbf{13.4}$$\mathbf{\pm 1}$
@@
Table 4: Ablation study on the number of Hydra’s heads ($G$) with MobileNet, CIFAR-10, $80\%$-DL, and IID:(80, 10).
Comparison with State-of-the-Art Methods. In Table 3, we evaluate Hydra (i.e., SFL+Hydra) using a random processing order (i.e., capturing a more generic scenario) and MobileNet. We compare it with other policies, i.e., SplitFedV1/FL (Thapa et al., 2022), SplitFedV3 (Madaan et al., 2022), Multihead (Chen et al., 2023), MergeSFL (Liao et al., 2024), and SplitNN Vepakomma et al. (2018). All of these methods have different workflows than SFL. Table 3 shows that Hydra significantly outperforms the state-of-the-art. Specifically, methods in the literature fail to improve on all metrics (indicating an unstable training), while Hydra consistently excels in all metrics (accuracy, PG, and BW). Importantly, Hydra outperforms the other multi-head approaches (i.e., SplitFedV3 and MultiHead). This highlights that Hydra’s aggregation of the heads is a key feature over the classic multi-head approaches. Moreover, MergeSFL achieves a global accuracy comparable to the one of Hydra. However, a significant drawback of MergeSFL lies in its implementation, which requires synchronization among clients and more frequent updates (due to micro-batches). Finally, among all the methods in the literature, SplitNN is the only one where the server processes clients’ data in a specific order (typically random). As a result, SplitNN has the worst (i.e., highest) performance gap among all methods due to forgetting.
Closing the Labels’ Performance Gap. To better understand Hydra’s performance, we dive into the per-position accuracy (i.e., the accuracy of the labels that are dominant in clients that are processed at a certain position, see Sec. 3.1). In Fig. 6, we see that Hydra effectively reduces the labels’ performance gap and stabilizes the training, even in a highly heterogeneous scenario, i.e., $80\%$-DL. In detail, unlike the baseline (SFL), there is no significant gap among the different positions. We note that the peaks and troughs in the subplot on the right are a result of the processing order under non-IID data. Finally, the global accuracy improves over SFL, i.e., $44\%$ instead of $28\%$, after 100 rounds.
Ablation Study on the Number of Heads $\mathbf{G}$. As discussed in Sec. 4, Hydra’s memory overhead can be adjusted by reducing the length of part-2b and/or $G$. For the former, we showed that a smaller part-2b performs better. Table 4 reports the performance of Hydra for different values of $G$ for CIFAR-10 and different processing orders of part-2a (cyclic with $\phi=10$ and random). For CIFAR-10, we define $G$ superclasses based on the semantic relationships among labels, similar to (Bai et al., 2021). For example, when $G=2$, we consider the two superclasses: animals and no-animals. We see that Hydra performs best when $G=L$, but still outperforms the baseline SFL in all metrics, even for a smaller number of heads. These results highlight an interesting trade-off between memory overhead and performance.
@@ 6 Conclusion and Discussion
Limitations and Future Work. Despite the absence of theoretical claims in this work, our empirical analysis of CF in SFL paves the way towards a theoretical/convergence analysis of SFL with respect to key parameters such as the cut layer and the processing order at the server. Such an analysis could also be the stepping stone towards establishing theoretical guarantees for Hydra. Furthermore, an interesting direction for future work is to incorporate client selection policies into the analysis of CF, but also in Hydra, similar to related work in FL, e.g., (Lee et al., 2022). Finally, our ablation study on the number of Hydra’s heads revealed that semantic relationships among labels may play an important role in CF and in SFL’s training in general. While this aspect was not the primary focus of this work, we believe that it should be studied in the future, as it has been done for continual learning (Ramasesh et al., 2021a).
Acknowledgments
This work was funded by the Horizon MSCA Postdoctoral Fellowship OPALS (grant agreement 101210495), the Spanish Ministry of Economic Affairs and Digital Transformation, the European Union-NextGenerationEU through the projects 6G-RIEMANN (TSI-063000-2021-147) and MAP-6G (TSI-063000-2021-63), and the Smart Networks and Services Joint Undertaking (SNS JU) under the European Union’s Horizon Europe research and innovation programme through the CONFIDENTIAL6G EU project (Grant Agreement No 101096435). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.